OpenStudy (anonymous):

A cylindrical vase has a diameter of 6 inches. At the bottom of the vase, there are 6 marbles, each of diameter 2 inches. The vase is filled with water up to a height of 12 inches. What is the volume of water in the vase? Answer 107π in3 102π in3 106π in3 108π in3

OpenStudy (anonymous):

Calculate the volume of the cylinder that is filled. Then subtract the volume of the marbles. You will then have the volume of the water.

OpenStudy (p0sitr0n):

\[v = pir^2h\] where V = colume pi = 3.1415 r = d/2 = 6/2 =3 h = 12 v =pi*9*12 v =108 pi same with marbles but those are spheres v = 4pir^3/3 6v = 24pi/3 6v = 8pi Vf = V cylinder - V marbles Vf = 100pi Vf =314,15

OpenStudy (anonymous):

oh wow thank you ! its 108 right?

OpenStudy (p0sitr0n):

100*pi = 314.15

OpenStudy (p0sitr0n):

you have to substract the marbles

OpenStudy (anonymous):

okay then i can up with 25.14 after subtracting the marbles?